OpenStudy (anonymous):

what is the solution of the following system -2x-y=1 -4x-2y=-1

OpenStudy (anonymous):

i got (3,8)

OpenStudy (skullpatrol):

Multiply both sides of the first equation by 2 :-)

OpenStudy (texaschic101):

lets see if (3,8) is correct... -2x - y = 1 -2(3) - 8 = 1 -6-8 = 1 - 14 = 1.....does that look true to you ?

OpenStudy (texaschic101):

so...do you prefer substitution or elimination ?

OpenStudy (texaschic101):

-2x - y = 1......-2x - 1 = y so sub in -2x-1 in for y in the other equation....then solve for x

OpenStudy (anonymous):

so no solutions?

OpenStudy (skullpatrol):

-2x - y = 1 2(-2x - y) = 2(1) -4x - 2y = 2

OpenStudy (skullpatrol):

Does 2 = -1? :-)

OpenStudy (anonymous):

no

OpenStudy (skullpatrol):

Therefore the solution set of the system is empty.

OpenStudy (skullpatrol):

Any questions?